# Misiurewicz points and subhyperbolicity in unicritical algebraic correspondences

### v1

@@ 4. Proof of Theorem A
Apply [ETDS22, Lemmas 2.1 and 2.2]. Since $a$ is Misiurewicz, we have only one bounded forward orbit of 0 under $\mathbf{f}_{a},$ which is strictly pre-periodic. There exist at most finitely many points of the postcritical set within any open ball $B(0,r).$
@@
We know that $\varphi:\mathbb{D}{\setminus}B\to\mathbb{C}{\setminus}R$ is a covering map, where $R=\varphi(B)$ is the set of ramified points. If $V$ is a sufficiently small open disk in $\mathbb{D}{\setminus}B$ then $\varphi(V)$ is also a small conformal disk in $\mathbb{C}{\setminus}R$ which is mapped under $\mathbf{f}_{a}^{-1}$ to $p$ disjoint conformal disks $D_{j}$ in $\mathbb{C}{\setminus}R,$ each of which evenly covered by $\varphi,$ that is, $\varphi$ maps every connected component of $\varphi^{-1}(D_{j})$ onto $D_{j}$ as a univalent map. Thus $\mathbf{h}(V)$ consists of disjoint conformal disks, each of which being the image of a univalent branch of $\mathbf{h}.$ The conditions that define a separable holomorphic multifunction (def. 3.3) are clearly implied by this property. ∎
@@
if $j=1$ and $z\in\Omega_{1}.$ Since $f_{1}$ coincide with $f_{2}$ on the intersection of their domains, (4.1) holds for $j=2$ and $z\in\Omega_{1}\cap\Omega_{2}.$ By the Identity Theorem, the corresponding equation (4.1) for $j=2$ actually holds for any $z$ in $\Omega_{2}.$ We may repeat this argument inductively. The conclusion: (4.1) holds for $z\in\Omega_{j}$, for every $j.$ In particular, for $j=n$ this implies $f_{n}(z)\in\varphi^{-1}\circ\mathbf{f}_{a}^{-1}\circ\varphi(z)$, for every $z\in\Omega_{n}.$ Hence $\mathfrak{h}=(f_{n},\gamma(1))$ is determined by a local branch of $\mathbf{h}.$
@@
Fix a point $z_{0}$ in $U.$ Any $z\in U$ is connected to $z_{0}$ by a curve $\gamma$ defined on $[0,1]$ with $\gamma(0)=z_{0}$ and $\gamma(1)=z.$ First, suppose that the analytic continuation $\tilde{\gamma}$ of any germ $\mathfrak{h}\in\Lambda_{\alpha}$ with $\pi(\mathfrak{h})=z_{0}$ along $\gamma$ always exists. Since $z$ is arbitrary and $\Lambda_{\alpha}$ is connected, it follows that $\tilde{\gamma}(1)$ is a germ in $\Lambda_{\alpha}$ whose projection is $z$, thereby proving that $\pi(\Lambda_{\alpha})=U.$ Now we will check that the continuation of any germ $\mathfrak{h}=(h,z_{0})$ along $\gamma$ exists. The local branch $h$ of $\mathbf{h}$ is defined on a neighborhood of $z_{0}$ and is given by the composition $g\circ f\circ\varphi,$ where $f$ is a local branch of $\mathbf{f}_{a}^{-1}$ at $\varphi(z_{0})$ and $g$ is a local inverse of $\varphi$ at $f(\varphi(z_{0})).$ Since $\gamma$ does not intersect $B$, its image $\gamma_{1}=\varphi\circ\gamma$ is contained in $\mathbb{C}{\setminus}R.$ Considering the global analytic multifunction $\mathbf{f}_{a}^{-1}:\mathbb{C}{\setminus}\{a\}\to\mathbb{C}{\setminus}\{0\}$, the continuation of $(f,\varphi(z_{0}))$ along $\gamma_{1}=\varphi\circ\gamma$ (see def. 3.1) yields a curve $\gamma_{2}$ in $\mathbb{C}{\setminus}R.$ As a consequence of the Lifting Theorem for covering spaces, the inverse $\varphi^{-1}:\mathbb{C}{\setminus}R\to\mathbb{D}{\setminus}B$ of the covering map is a global analytic multifunction and $(g,f(\varphi(z_{0}))$ is a germ of $\varphi^{-1}$; the continuation of this germ along $\gamma_{2}$ yields a curve $\gamma_{3}.$ Notice that $\gamma_{3}(t)$ is in $\mathbf{h}(\gamma(t))$, for every $t.$ Taking into account Lemma 4.2 and Theorem 3.2, the final curve $\gamma_{3}$ is given by an analytic continuation of $\mathfrak{h}\in\mathfrak{S}(\mathbf{h})$ along $\gamma,$ as desired. ∎
@@
We must define $\mathbf{h}_{\alpha}$ explicitly. By definition, if $z\in U,$ then $w$ belongs to $\mathbf{h}_{\alpha}(z)$ whenever $w=h(z)$, for some germ of analytic function $\mathfrak{h}=(h,z)$ in $\Lambda_{\alpha}$ with $\pi(\mathfrak{h})=z.$ In this case, it should be noticed that $(h,\zeta)$ is in $\Lambda_{\alpha}$ for every $\zeta$ in the domain of $h,$ for then $\Lambda_{\alpha}$ is a connected component of the sheaf $\mathfrak{S}_{U}.$ This shows that every $h$ used to define a value of $\mathbf{h}_{\alpha}(z)$ can also be used to define $h(\zeta)$ as a value of $\mathbf{h}_{\alpha}(\zeta),$ for every $\zeta$ in the domain of $h.$ Hence $h$ is a branch of the multifunction $\mathbf{h}_{\alpha}.$ It follows that $\mathbf{h}_{\alpha}$ is a holomorphic multifunction.
@@
For the converse inclusion $\mathfrak{S}(\mathbf{h}_{\alpha})\subset\Lambda_{\alpha}$, assume $\mathfrak{h}$ is a germ in $\mathfrak{S}(\mathbf{h}_{\alpha}).$ Then $\mathfrak{h}=(h,z_{0})$, for some branch $h$ of $\mathbf{h}$ and some $z_{0}$ in the domain of $h.$ Using the definition of $\mathbf{h}_{\alpha}$, for every $z$ in the domain of $h$ we find a holomorphic map $g_{z}$ locally defined at $z$ such that $(g_{z},z)$ is germ in $\Lambda_{\alpha}$ and $h(z)=g_{z}(z).$ By Lemma 4.3, $\Lambda_{\alpha}$ is a connected component of the sheaf, and
@@
for every $\zeta$ in the domain of $g_{z}.$ This means that $g_{z}$ is a branch of $\mathbf{h}$ when restricted to some small neighborhood of $\zeta$, for every $\zeta$ in the domain of $g_{z}.$ We conclude that $g_{z}$ is a branch of $\mathbf{h}.$ Since $(z)=g_{z}(z), hh$ is also a branch of $\mathbf{h}.$ By step $(a),$ $\mathbf{h}$ is separable with $E(\mathbf{h})=\emptyset.$ Both $g_{z_{0}}$ and $h$ are branches of $\mathbf{h}$ locally defined at $z_{0}.$ According to Definition 3.3, for some neighborhood $V$ of $z_{0}$, either $h(V)$ and $g_{z_{0}}(V)$ are disjoint, or else $h=g_{z_{0}}$ when restricted to $V.$ But $h(z_{0})=g_{z_{0}}(z_{0})$. Therefore,
@@
For the existence of the continuation of $\mathfrak{h}=(h,z_{0})$ along $\gamma$ we first notice that $\gamma$ projects onto $\varphi\circ\gamma.$ Since $\mathbf{f}_{a}^{-1}$ is an algebraic multifunction on $\mathbb{C}{\setminus}\{a\}$, any germ of $\mathbf{f}_{a}^{-1}$ at the first point of $\varphi\circ\gamma$ can be continued along this curve. The continuation yields a curve $\zeta$ in the plane avoiding the set of ramified points $R$ such that $\zeta(t)\in\mathbf{f}_{a}^{-1}(\varphi\circ\gamma(t))$, for every $t,$ with $\zeta(0)=\zeta_{0},$ where $\zeta_{0}=\varphi(w_{0})$ and $w_{0}=h(z_{0}).$ Since $\varphi:\mathbb{D}{\setminus}B\to\mathbb{C}{\setminus}R$ is a covering map, by the lifting theorem there exists a unique curve $\eta$ in the disk which projects onto $\zeta=\varphi\circ\eta.$ Now $\eta(t)\in\mathbf{h}_{\alpha}(\gamma(t))$ for every $t$, so that by Theorem 3.2 $\eta$ is given by an analytic continuation $\tilde{\gamma}$ of $\mathfrak{h}$ along $\gamma.$ In particular, the continuation exists. ∎
@@
There are infinitely many points $y$ in $\mathbf{h}_{\alpha}(x)$. However, in the Lemma 4.7 we will make the choice of $b$ and $y$ unique, depending only on $x$ and $\mathfrak{h}=(h,x+r).$
By step Lemma 4.5, if $\zeta:[0,1]\to U\cup\{x\}$ is a curve starting at $x+r$ and terminating at $x,$ then the analytic continuation of $\mathfrak{h}$ along the restriction of $\zeta$ to any closed subinterval $0,t]\subset[0,1)$ exists. If we take $t\to 1$ we conclude that there exists an extension of the domain of $h$ to an open set $V\subset U$ with a sequence $x_{k}\in V$ converging to $x$, such that $V$ has diameter $r+\epsilon$, for $\epsilon>0$ arbitrarily small. (Alternatively, we may apply Corollary 3.1 to reach the same conclusion, requiring $V$ to be simply connected).
@@
The local branch $h$ is given by a composition $\psi\circ f^{-1}\circ\varphi$, where $f^{-1}$ is univalent branch of $\mathbf{f}_{a}^{-1}$, $\psi$ is a univalent branch of $\varphi^{-1}$. Since $\varphi$ is continuous, $\varphi(x_{k})$ converges to $a_{j}=\varphi(x).$ The sequence $f^{-1}\circ\varphi(x_{k})$ also converges to a point $b$ which is independent of $\{x_{k}\}$ and $V$: if $a=a_{j}$ then take $b=0$; otherwise, $f^{-1}$ can be extended to a domain that includes $a_{j}$, for then $a_{j}$ is in the closure of the domain of $f^{-1}$. In this case, $b=f^{-1}(a_{j}).$
Since $f^{-1}$ is branch of $\mathbf{f}_{a}^{-1}$, which is a continuous set-valued function in the Hausdorff topology, if the diameter of $V$ is small, then $f^{-1}\circ\varphi(V)$ is also small. Hence we may assume that $f^{-1}\circ\varphi(V)$ is contained in a connected $W$, with $b\in W$, such that $W$ is evenly covered by $\varphi$, that is, each connected component of $\varphi^{-1}(W)$ projects onto $W$ by means of a proper map. By applying $\psi$, which is a local inverse of $\varphi$, we conclude that $h(V)$ is contained in one of such components, say $C_{0}.$ If $\nu(b)=1$, by reducing $V$ (and consequently $W$), we may assume that $\varphi:C_{0}\to W$ is bi-holomorphic. It is clear that $h(x_{k})$ converges to $(\varphi|_{C_{0}})^{-1}(b)$, which is independent of $V$ and $\{x_{k}\}.$
@@
Since $r$ is small and $a_{j}\neq a,$ the univalent map $f^{-1}$ can be extended to a connected neighborhood of $a_{j}$ that includes $\{\tilde{\gamma}\}$. The image curve $\tilde{\eta}=f^{-1}\circ\tilde{\gamma}$ also has winding number $n(\tilde{\eta},b)$ equal to $p.$ Now $b=a_{k}$ is a ramified point with index $\nu(a_{k})=p$, and the local degree of $\varphi|_{C_{0}}$ at $y$ is precisely $p,$ for then $\varphi|_{C_{0}}$ is conjugate to $z^{p}$, up to a translation, as described in the proof of $(e1).$ (The following term yields has a rigorous meaning, see Definition 3.1). Hence the analytic continuation of the germ
@@
which equals $(1,p,p,1).$ The case $(\nu(a_{j}),\nu(b))=(q,q)$ is impossible, for then $a_{j}=0$ and $b=0,$ and the critical point could not be strictly pre-periodic (Misiurewicz). Case $(q,p)$ is also impossible, by a similar reasoning. In all other cases, we can obtain different sequences of curves and winding numbers, which we indicate using the same notation of (4.4). In case $(p,q)$, for example, we only have to take into account that in the equation $(w-c)^{q}=z^{p}$ which defines $\mathbf{f}_{a}$, the continuation of a curve $\zeta$ with winding number $n(\zeta,0)=q$ in the $z$-plane yields a curve $\tilde{\zeta}$ with winding number $n(\tilde{\zeta},a)=p$ in the $w$-plane. We summarize as follows:
@@
The values of the extended $\mathbf{g}_{\alpha}$ at every singularity $x=x_{j}$ are determined by Theorem 3.4. According to this result, every branch of $\mathbf{h}_{\alpha}$ locally defined at $x_{j}+r$ can be extended to a punctured neighborhood of $x_{j}$, with a removable singularity at $x_{j}$, thus satisfying the conditions of step $(e1),$ with $y=y_{j}.$ Hence the set $\mathbf{g}_{\alpha}(x_{j})$ is given by all
@@
In a second case, the only branch points in $\gamma$ are the starting and terminal points. It is easy to determine a continuation $\tilde{\gamma}(t)$ for $t$ sufficiently close to $0.$ Since $\gamma$ contains no branch points on $(0,1)$, there exists a unique extension of $\tilde{\gamma}$ to a continuation along $\gamma|_{[0,t]}$ for $t$ arbitrarily close to $1.$ To define the value of the continuation near $1$ we apply Theorem 3.4: if $t^{*}$ is sufficiently close to $1$, then there exists a unique branch $F$ of $\mathbf{g}_{\alpha}$ locally defined at $\gamma(1)$ such that $(F,\gamma(t^{*}))=\tilde{\gamma}(t^{*}).$ Since any two analytic continuations are either identical or else differ for every $t,$ we have $(F,\gamma(t))=\tilde{\gamma}(t)$ as long as $t<1$ and $\gamma(t)$ is in the domain of $F.$ If we set $\tilde{\gamma}(1)=(F,\gamma(1))$, the result is a continuous curve $\tilde{\gamma}$ in the sheaf space; in other words, an analytic continuation along $\gamma.$

### v2

@@ 4. Proof of Theorem A
Apply [14, Lemmas 2.1 and 2.2]. Since $a$ is Misiurewicz, we have only one bounded forward orbit of 0 under $\mathbf{f}_{a},$ which is strictly preperiodic. There exist at most finitely many points of the postcritical set within any open ball $B(0,r).$
@@
We know that $\varphi:\mathbb{D}{\setminus}B\to\mathbb{C}{\setminus}R$ is a covering map, where $R=\varphi(B)$ is the set of ramified points. If $V$ is a sufficiently small open disk in $\mathbb{D}{\setminus}B$ then $\varphi(V)$ is also a small conformal disk in $\mathbb{C}{\setminus}R$ which is mapped under $\mathbf{f}_{a}^{-1}$ to $p$ disjoint conformal disks $D_{j}$ in $\mathbb{C}{\setminus}R,$ each of which evenly is covered by $\varphi,$ that is, $\varphi$ maps every connected component of $\varphi^{-1}(D_{j})$ onto $D_{j}$ as a univalent map. Thus $\mathbf{h}(V)$ consists of disjoint conformal disks, each of which being the image of a univalent branch of $\mathbf{h}.$ The conditions that define a separable holomorphic multifunction (def. 3.3) are clearly implied by this property. ∎
@@
if $j=1$ and $z\in\Omega_{1}.$ Since $f_{1}$ coincides with $f_{2}$ on the intersection of their domains, (4.1) holds for $j=2$ and $z\in\Omega_{1}\cap\Omega_{2}.$ By the Identity Theorem, the corresponding equation (4.1) for $j=2$ actually holds for any $z$ in $\Omega_{2}.$ We may repeat this argument inductively. The conclusion: (4.1) holds for $z\in\Omega_{j}$, for every $j.$ In particular, for $j=n$ this implies $f_{n}(z)\in\varphi^{-1}\circ\mathbf{f}_{a}^{-1}\circ\varphi(z)$, for every $z\in\Omega_{n}.$ Hence $\mathfrak{h}=(f_{n},\gamma(1))$ is determined by a local branch of $\mathbf{h}.$
@@
Fix a point $z_{0}$ in $U.$ Any $z\in U$ is connected to $z_{0}$ by a curve $\gamma$ defined on $[0,1]$ with $\gamma(0)=z_{0}$ and $\gamma(1)=z.$ First, suppose that the analytic continuation $\tilde{\gamma}$ of any germ $\mathfrak{h}\in\Lambda_{\alpha}$ with $\pi(\mathfrak{h})=z_{0}$ along $\gamma$ always exists. Since $z$ is arbitrary and $\Lambda_{\alpha}$ is connected, it follows that $\tilde{\gamma}(1)$ is a germ in $\Lambda_{\alpha}$ whose projection is $z$, thereby proving that $\pi(\Lambda_{\alpha})=U.$ Now we will check that the continuation of any germ $\mathfrak{h}=(h,z_{0})$ along $\gamma$ exists. The local branch $h$ of $\mathbf{h}$ is defined on a neighborhood of $z_{0}$ and is given by the composition $g\circ f\circ\varphi,$ where $f$ is a local branch of $\mathbf{f}_{a}^{-1}$ at $\varphi(z_{0})$ and $g$ is a local inverse of $\varphi$ at $f(\varphi(z_{0})).$ Since $\gamma$ does not intersect $B$, its image $\gamma_{1}=\varphi\circ\gamma$ is contained in $\mathbb{C}{\setminus}R.$ Considering the global analytic multifunction $\mathbf{f}_{a}^{-1}:\mathbb{C}{\setminus}\{a\}\to\mathbb{C}{\setminus}\{0\}$, the continuation of $(f,\varphi(z_{0}))$ along $\gamma_{1}=\varphi\circ\gamma$ (see def. 3.1) yields a curve $\gamma_{2}$ in $\mathbb{C}{\setminus}R.$ As a consequence of the Lifting Theorem for covering spaces, the inverse $\varphi^{-1}:\mathbb{C}{\setminus}R\to\mathbb{D}{\setminus}B$ of the covering map is a global analytic multifunction and $(g,f(\varphi(z_{0})))$ is a germ of $\varphi^{-1}$; the continuation of this germ along $\gamma_{2}$ yields a curve $\gamma_{3}.$ Notice that $\gamma_{3}(t)$ is in $\mathbf{h}(\gamma(t))$, for every $t.$ Taking into account Lemma 4.2 and Theorem 3.2, the final curve $\gamma_{3}$ is given by an analytic continuation of $\mathfrak{h}\in\mathfrak{S}(\mathbf{h})$ along $\gamma,$ as desired. ∎
@@
We must define $\mathbf{h}_{\alpha}$ explicitly. By definition, if $z\in U,$ then $w$ belongs to $\mathbf{h}_{\alpha}(z)$ whenever $w=h(z)$, for some germ $\mathfrak{h}=(h,z)$ in $\Lambda_{\alpha}$ with $\pi(\mathfrak{h})=z.$ In this case, it should be noticed that $(h,\zeta)$ is in $\Lambda_{\alpha}$ for every $\zeta$ in the domain of $h,$ for then $\Lambda_{\alpha}$ is a connected component of the sheaf $\mathfrak{S}_{U}.$ This shows that every $h$ used to define a value of $\mathbf{h}_{\alpha}(z)$ can also be used to define $h(\zeta)$ as a value of $\mathbf{h}_{\alpha}(\zeta),$ for every $\zeta$ in the domain of $h.$ Hence $h$ is a branch of the multifunction $\mathbf{h}_{\alpha}.$ It follows that $\mathbf{h}_{\alpha}$ is a holomorphic multifunction.
@@
For the converse inclusion $\mathfrak{S}(\mathbf{h}_{\alpha})\subset\Lambda_{\alpha}$, assume $\mathfrak{h}$ is a germ in $\mathfrak{S}(\mathbf{h}_{\alpha}).$ Then $\mathfrak{h}=(h,z_{0})$, for some branch $h$ of $\mathbf{h}$ and some $z_{0}$ in the domain of $h.$ Using the definition of $\mathbf{h}_{\alpha}$, for every $z$ in the domain of $h$ we find a holomorphic map $g_{z}$ locally defined at $z$ such that $(g_{z},z)$ is a germ in $\Lambda_{\alpha}$ and $h(z)=g_{z}(z).$ By Lemma 4.3, $\Lambda_{\alpha}$ is a connected component of the sheaf, and
@@
for every $\zeta$ in the domain of $g_{z}.$ This means that $g_{z}$ is a branch of $\mathbf{h}$ when restricted to some small neighborhood of $\zeta$, for every $\zeta$ in the domain of $g_{z}.$ We conclude that $g_{z}$ is a branch of $\mathbf{h}.$ Since $(z)=g_{z}(z), hh$ is also a branch of $\mathbf{h}.$ By Lemma 4.2, $\mathbf{h}$ is separable with $E(\mathbf{h})=\emptyset.$ Both $g_{z_{0}}$ and $h$ are branches of $\mathbf{h}$ locally defined at $z_{0}.$ According to Definition 3.3, for some neighborhood $V$ of $z_{0}$, either $h(V)$ and $g_{z_{0}}(V)$ are disjoint, or else $h=g_{z_{0}}$ when restricted to $V.$ But $h(z_{0})=g_{z_{0}}(z_{0})$. Therefore,
@@
For the existence of the continuation of $\mathfrak{h}=(h,z_{0})$ along $\gamma$ we first notice that $\gamma$ projects onto $\varphi\circ\gamma.$ Since $\mathbf{f}_{a}^{-1}$ is an algebraic multifunction on $\mathbb{C}{\setminus}\{a\}$, any germ of $\mathbf{f}_{a}^{-1}$ at the first point of $\varphi\circ\gamma$ can be continued along this curve. The continuation yields a curve $\zeta$ in the plane avoiding the set of ramified points $R$ such that $\zeta(t)\in\mathbf{f}_{a}^{-1}(\varphi\circ\gamma(t))$, for every $t,$ with $\zeta(0)=\zeta_{0},$ where $\zeta_{0}=\varphi(w_{0})$ and $w_{0}=h(z_{0}).$ Since $\varphi:\mathbb{D}{\setminus}B\to\mathbb{C}{\setminus}R$ is a covering map, by the lifting theorem, there exists a unique curve $\eta$ in the disk with $\eta(0)=w_{0}$ such that $\varphi\circ\eta=\zeta$. Now $\eta(t)\in\mathbf{h}_{\alpha}(\gamma(t))$ for every $t$, so that by Theorem 3.2 $\eta$ is given by an analytic continuation $\tilde{\gamma}$ of $\mathfrak{h}$ along $\gamma.$ In particular, the continuation exists. ∎
@@
In Lemma 4.7 we will make the choice of $b$ and $y$ unique, depending only on $x$ and $\mathfrak{h}=(h,x+r).$
By Lemma 4.5, if $\zeta:[0,1]\to U\cup\{x\}$ is a curve starting at $x+r$ and terminating at $x,$ then the analytic continuation of $\mathfrak{h}$ along the restriction of $\zeta$ to any closed subinterval $0,t]\subset[0,1)$ exists. If we take $t\to 1$ we conclude that there exists an extension of the domain of $h$ to an open set $V\subset U$ with a sequence $x_{k}\in V$ converging to $x$, such that $V$ has diameter $r+\epsilon$, for $\epsilon>0$ arbitrarily small. (Alternatively, we may apply Corollary 3.1 to reach the same conclusion, requiring $V$ to be simply connected).
@@
The local branch $h$ is given by a composition $\psi\circ f^{-1}\circ\varphi$, where $f^{-1}$ is a univalent branch of $\mathbf{f}_{a}^{-1}$, $\psi$ is a univalent branch of $\varphi^{-1}$. Since $\varphi$ is continuous, $\varphi(x_{k})$ converges to $a_{j}=\varphi(x).$ The sequence $f^{-1}\circ\varphi(x_{k})$ also converges to a point $b$ which is independent of $\{x_{k}\}$ and $V$: if $a=a_{j}$ then take $b=0$; otherwise, $f^{-1}$ can be extended to a domain that includes $a_{j}$, for then $a_{j}$ is in the closure of the domain of $f^{-1}$. In this case, $b=f^{-1}(a_{j}).$
Since $f^{-1}$ is branch of $\mathbf{f}_{a}^{-1}$, which is a continuous set-valued function in the Hausdorff topology, if the diameter of $V$ is small, then $f^{-1}\circ\varphi(V)$ is also small. Hence we may assume that $f^{-1}\circ\varphi(V)$ is contained in a connected $W$, with $b\in W$, such that $W$ is evenly covered by $\varphi$, that is, each connected component of $\varphi^{-1}(W)$ projects onto $W$ by means of a proper map. By applying $\psi$, which is a local inverse of $\varphi$, we conclude that $h(V)$ is contained in one of these components, say $C_{0}.$ If $\nu(b)=1$, by reducing $V$ (and consequently $W$), we may assume that $\varphi:C_{0}\to W$ is bi-holomorphic. It is clear that $h(x_{k})$ converges to $(\varphi|_{C_{0}})^{-1}(b)$, which is independent of $V$ and $\{x_{k}\}.$
@@
Since $r$ is small and $a_{j}\neq a,$ the univalent map $f^{-1}$ can be extended to a connected neighborhood of $a_{j}$ that includes $\{\tilde{\gamma}\}$. The image curve $\tilde{\eta}=f^{-1}\circ\tilde{\gamma}$ also has winding number $n(\tilde{\eta},b)$ equal to $p.$ Now $b=a_{k}$ is a ramified point with index $\nu(a_{k})=p$, and the local degree of $\varphi|_{C_{0}}$ at $y$ is precisely $p,$ for then $\varphi|_{C_{0}}$ is conjugate to $z^{p}$, up to a translation, as described in the proof of Lemma 4.7. (The following term yields has a rigorous meaning, see Definition 3.1). Hence the analytic continuation of the germ
@@
which equals $(1,p,p,1).$ The case $(\nu(a_{j}),\nu(b))=(q,q)$ is impossible, for then $a_{j}=0$ and $b=0,$ and the critical point could not be strictly preperiodic (Misiurewicz). Case $(q,p)$ is also impossible, by a similar reasoning. In all other cases, we can obtain different sequences of curves and winding numbers, which we indicate using the same notation from (4.4). In case $(p,q)$, for example, we only have to take into account that in the equation $(w-c)^{q}=z^{p}$ which defines $\mathbf{f}_{a}$, the continuation of a curve $\zeta$ with winding number $n(\zeta,0)=q$ in the $z$-plane yields a curve $\tilde{\zeta}$ with winding number $n(\tilde{\zeta},a)=p$ in the $w$-plane. We summarize as follows:
@@
The values of the extended $\mathbf{g}_{\alpha}$ at every singularity $x=x_{j}$ are determined by Theorem 3.4. According to this result, every branch of $\mathbf{h}_{\alpha}$ locally defined at $x_{j}+r$ can be extended to a punctured neighborhood of $x_{j}$, with a removable singularity at $x_{j}$, thus satisfying the conditions of Lemma 4.7 with $y=y_{j}.$ Hence the set $\mathbf{g}_{\alpha}(x_{j})$ is given by all
@@
In a second case, the only branch points in $\gamma$ are the starting and terminal points. It is easy to determine a continuation $\tilde{\gamma}(t)$ for $t$ sufficiently close to $0.$ Since $\gamma$ contains no branch points on $(0,1)$, there exists a unique extension of $\tilde{\gamma}$ to a continuation along $\gamma|_{[0,t]}$ for $t$ arbitrarily close to $1.$ To define the value of the continuation near $1$ we apply Theorem 3.4: if $t^{*}$ is sufficiently close to $1$, then there exists a unique branch $F$ of $\mathbf{g}_{\alpha}$ locally defined at $\gamma(1)$ such that $(F,\gamma(t^{*}))=\tilde{\gamma}(t^{*}).$ Since any two analytic continuations are either identical or else differ for every $t,$ we have $(F,\gamma(t))=\tilde{\gamma}(t)$ as long as $t<1$ and $\gamma(t)$ is in the domain of $F.$ If we set $\tilde{\gamma}(1)=(F,\gamma(1))$, the result is a continuous curve $\tilde{\gamma}$ in the sheaf; in other words, an analytic continuation along $\gamma.$